# Learning-based Multi-View Stereo: A Survey

### v2

@@ 1 Introduction
One fundamental technique in image-based 3D reconstruction is Multi-View Stereo (MVS). Given a set of calibrated images, MVS aims to reconstruct dense 3D geometry for an observed scene. Based on the scene representations, traditional MVS methods can be divided into three categories: volumetric, point cloud, and depth map. Volumetric methods [9, 10, 11, 12] discretize the 3D space into voxels and label each as inside or outside of the surface. They are limited to scenes of small scale due to the large memory consumption. Point cloud methods [13, 14] operate directly on 3D points and often employ propagation to gradually densify the reconstruction. As point cloud propagation occurs sequentially, these methods are difficult to parallelize, leading to longer processing times. [15]. In addition, the irregularity and large size of point cloud are also not very suitable for deep learning, especially in large-scale scenes. In contrast, methods relying on depth maps [16, 17, 18, 19, 20] use patch matching with photometric consistency to estimate the depth maps for individual images. Subsequently, these depth maps are fused into a dense representation, e.g., point cloud or mesh, as a post-processing step. Such design decouples the reconstruction problem into per-view depth estimation and depth fusion, which explicitly improves the flexibility and scalability. Although MVS has been studied extensively for several decades, traditional MVS methods rely on hand-crafted matching metrics and thus encounter challenges in handling various conditions, e.g., illumination changes, low-textured areas, and non-Lambertian surfaces [21, 22, 23, 24].
@@
In this survey, we categorize existing learning-based MVS methods based on their characteristics as follows: depth map-based, voxel-based, NeRF-based, 3D Gaussian Splatting-based and large feed-forward methods. Voxel-based methods estimate the geometry with volumetric representation and implicit function, e.g., Signed Distance Functions (SDF). They are limited to small-scale scenes due to the high memory consumption. NeRF and 3D Gaussian Splatting-based methods adapt NeRF [27] and 3D Gaussian Splatting [28], which are originally used for novel view synthesis, to extract surface from the implicit field or point cloud. They typically need to optimize the geometry for each new scene, which needs lots of run-time and memory consumption. Large feed-forward methods typically use large transformer models to directly learn the 3D representation from given images. They require massive computation because of the huge network, and are mainly limited to object-level scenes [29]. Depth map-based methods utilize deep learning in depth estimation and then fuse depth maps with traditional fusion algorithm. Comparatively, depth map-based methods [15, 25] are the main family of learning-based methods since they inherit the advantages from those traditional methods [17, 18, 19, 20] by decoupling 3D reconstruction into depth estimation and depth fusion. Therefore, we focus more on the depth map-based methods and discuss them in details.
For clarity, we further categorize depth map-based methods into online and offline methods, shown in Fig. 1. Specifically, online MVS methods [25, 30] usually perform multi-view depth estimation with a video sequence of low-resolution images, e.g., ScanNet [26].These methods typically rely on low-resolution inputs and lightweight network structures to ensure fast inference, prioritizing speed and simplicity. The main goal of online MVS is to deliver real-time reconstruction for time-sensitive applications such as augmented reality and live video processing. In contrast, offline MVS methods [15, 31] prioritize better quality reconstruction at the cost of computation. These methods excel in multi-view depth estimation and point cloud fusion using high-resolution image sets, e.g., DTU [21], Tanks and Temples [22] and ETH3D [23]. By operating on higher-resolution inputs and employing complex network, they thoroughly analyze the multi-view information in the images. The emphasis on achieving high accuracy and capturing intricate details often comes at the cost of requiring considerable computational resources and time. As a result, offline MVS is frequently employed in applications demanding precise and photorealistic scene representations, including 3D modeling and archaeological reconstruction.
In summary, our survey covers the most recent literature on learning-based MVS methods, including four main families: depth map-based, voxel-based, NeRF-based, 3D Gaussian Splatting-based and large feed-forward methods. We provide a comprehensive review and insights on different aspects, including the pipelines and algorithmic intricacies. Moreover, we summarize the performance of the reviewed methods on different benchmarks, and discuss the potential future directions for deep learning-based MVS methods.
@@ 2 Preliminaries
Depth map-based MVS, including most traditional and learning-based methods, typically consists of several components: camera calibration, view selection, multi-view depth estimation and depth fusion. In this section, we introduce these components to provide readers a clear picture of the MVS. Note that camera calibration and view selection are components for other learning-based methods as well. Sec. 2.1 introduces camera calibration with Structure from Motion (SfM) or SLAM. Sec. 2.2 discusses how to select neighboring views to reconstruct the geometry. Sec. 2.3 explains how to build cost volumes in learning-based MVS methods with plane sweep [32]. Sec. 2.4 introduces the typical depth fusion strategies after depth estimation. Sec. 2.5 lists common datasets and benchmarks for MVS and Sec. 2.6 summarizes the common evaluation metrics.
{forest}
for tree= forked edges, grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=middle, font=, rectangle, draw=black, rounded corners,align=left, minimum width=2.5em, minimum height=1.2em, s sep=6pt, inner xsep=3pt, inner ysep=1pt, , where level=1text width=4.5em, where level=2text width=8em,font=, where level=3font=, where level=4font=, where level=5font=, [MVS [Traditional Methods,text width=10 em [COLMAP [33], P-MVS [14], Gipuma [18]…,leaf,text width=16.3 em] ] [Datasets & Benchmarks, text width=10 em [ScanNet [26], DTU [21], Tanks and Temples [22]…, leaf,text width=16.3 em] ] [Learning-based Methods
with Depth Estimation,text width=10 em [Pipeline,text width=6 em [Camera Calibration, View Selection, Multi-View Depth Estimation, Depth Fusion,leaf,text width=27.4 em] ] [Supervised,text width=6em [Online Methods [MVDepthNet [25], DeepvideoMVS [30], SimpleRecon [34]…,leaf,text width=20.2 em] ] [Offline Methods [Direct 3D CNN,text width=5.4 em [MVSNet [15], CIDER [35]…,leaf,text width=13 em] ] [RNN,text width=5.4 em [R-MVSNet [36], $D^{2}$-RMVSNet [37]…,leaf,text width=13 em] ] [Coarse-to-fine,text width=5.4 em [CasMVSNet [31], UCSNet [38]…,leaf,text width=13 em] ] [Iterative Update,text width=5.4 em [PatchmatchNet [39], IterMVS [40]…,leaf,text width=13 em] ] ] ] [Unsupervised,text width=6em [End-to-end,text width=5 em [ JDACS [41], RC-MVSNet [42]…,leaf,text width=10.5 em] ] [Multi-stage,text width=5 em [U-MVSNet [43], KD-MVS [44]…,leaf,text width=10.5 em] ] ] [Semi-Supervised,text width=6em [ SGT-MVSNet [45],leaf,text width=6.2 em] ] ] [Learning-based Methods
without Depth Estimation,text width=10em [Voxel-based,text width=10 em [Atlas [46], NeuralRecon [47]…, leaf,text width=13 em] ] [NeRF-based,text width=10 em [VolSDF [48], NeuS [49], Geo-NeuS [50]…, leaf,text width=13 em] ] [3D Gaussian Splatting-based,text width=10 em [ SuGaR [51], 2DGS [52], PGSR [53]…, leaf,text width=13 em] ] [Large feed-forward,text width=10 em [LRM [54], DUSt3R [55], MASt3R [56]…, leaf,text width=13 em] ] ] [Future Research Directions, text width=10 em [Dataset & Benchmarks, View Selection, Depth Fusion, Features, Run-time & Memory Efficiency, Prior Assistance, leaf,text width=37.5 em] ] ]
@@ 2.1 Camera Calibration
[…]
@@ 2.2 View Selection
First, for most online MVS depth estimation methods [25, 60, 30], a frame is selected as a keyframe when its pose has sufficient difference compared with the previous keyframe. Then each keyframe adopts several previous keyframes to estimate depth. GP-MVS [60] proposes a heuristic pose-distance measure as:
$\text{disc}(\mathbf{T}_{ij})=\sqrt{||\mathbf{t}_{ij}||^{2}+\frac{2}{3}\text{tr}(\mathbbm{I}-\mathbf{R}_{ij})},$ (1)
where $\mathbf{T}_{ij}=[\mathbf{R}_{ij}|\mathbf{t}_{ij}]$ is the relative transformation between view $i$ and $j$. This strategy is used by many following methods [30, 34].
Second, for most offline MVS methods [15, 36, 31], view selection is done with the sparse point cloud obtained by Structure-from-Motion [33, 58]. For a reference view $i$, MVSNet [15] computes a score $s(i,j)=\sum_{\mathbf{P}}\eta(\theta_{ij}(\mathbf{P}))$ for the neighboring view $j$, where $\mathbf{P}$ is a 3D point observed by both view $i$ and $j$. $\theta_{ij}(\mathbf{P})=(180/\pi)\arccos{((\mathbf{c}_{i})-\mathbf{P})\cdot((\mathbf{c}_{j})-\mathbf{P}))}$ represents the baseline angle for $\mathbf{P}$ and $\mathbf{c}_{i},\mathbf{c}_{j}$ are the camera centers. $\eta(\cdot)$ is a piece-wise Gaussian function [61] to favor a certain baseline angle $\theta_{0}$:
$\eta(\theta)=\left\{\begin{aligned} \exp{(-\frac{(\theta-\theta_{0})^{2}}{2\sigma_{1}^{2}})},\theta\leq\theta_{0}\\ \exp{(-\frac{(\theta-\theta_{0})^{2}}{2\sigma_{2}^{2}})},\theta>\theta_{0}\end{aligned}\right.$ (2)
where $\theta_{0}$, $\sigma_{0}$, $\sigma_{1}$ are hyper-parameters. Then the view selection is done by choosing neighboring views with highest scores. Almost all the following offline MVS methods [36, 31, 39] use the same strategy.
@@ 2.3 Multi-view Depth Estimation with Plane Sweep
To form a structured data format that is more suitable for convolution operations, most learning-based MVS methods rely on the plane sweep algorithm [32] to calculate matching costs. The plane sweep algorithm discretizes the depth space with a set of fronto-parallel planes along the depth direction. This practice is deeply inspired by learning-based binocular stereo methods [62, 63, 64], which assess matching costs for a set of disparity hypotheses and subsequently estimate the disparity.
@@
In practice, we divide the depth range, which is either manually set [25] or estimated by SfM [15], into discrete samples and assign hypotheses at these values. To map coordinates at depth hypothesis $d$, homography transformation [15] $\mathbf{p}_{i}(d)\sim\mathbf{H}_{i}(d)\cdot\mathbf{p}$, is applied, where $\mathbf{p}_{i}(d)$ is the corresponding pixel in the $i$-th source view for pixel $\mathbf{p}$ of the reference view. The homography $\mathbf{H}_{i}(d)$ between the $i$-th source view and the reference view 0 can be computed as:
@@
where $\mathbf{K}_{0}$, $\mathbf{K}_{i}$ denote camera intrinsics, $[\mathbf{R}_{0}|\mathbf{t}_{0}]$, and $[\mathbf{R}_{i}|\mathbf{t}_{i}]$ denote camera extrinsics, $\mathbf{n}$ denote the principle axis of the reference view. Equivalently, we can also project a reference pixel into source views with depth hypothesis [35, 39, 40]. We compute $\mathbf{p}_{i}(d)$ in the $i$-th source view for pixel $\mathbf{p}$ in the reference and depth hypothesis $d$ as follows:
@@
The warped source feature maps are then obtained via differentiable interpolation. To better elaborate the process of plane sweep, we include a toy example in the supplementary. The photometric similarity (or matching cost) is measured in a one-to-many manner between reference and warped source features for the following depth estimation, which will be discussed in Sec. 3.2. By performing plane sweep, we obtain a cost volume with a regular spatial shape, making it easy for CNNs to process in parallel.
@@ 2.4 Depth Fusion
For depth map-based MVS, after estimating all the depth maps, we need to fuse them into a dense 3D representation, e.g., point cloud or mesh. Online MVS methods [30, 34] usually adopt TSDF (Truncated Signed Distance Function) fusion [65, 66] to fuse the depth maps into a TSDF volume and then use Marching Cube [67] to extract the mesh. However, there usually exist outliers in the depth maps, which may reduce the reconstruction accuracy. To overcome this problem and improve the accuracy, offline MVS methods [15, 31, 39] typically filter the depth maps before fusing into a point cloud, which is motivated by Galliani et al. [18]. There are two main filtering steps: photometric consistency filtering and geometric consistency filtering [15]. For more details, please refer to the supplementary.
@@ 2.5 Datasets and Benchmarks
Along with this survey, we also include a brief summary of commonly used public MVS datasets and benchmarks for training and evaluation. For more details, please refer to supplementary.

### v3

@@ 1 Introduction
One fundamental technique in image-based 3D reconstruction is Multi-View Stereo (MVS). Given a set of calibrated images, MVS aims to reconstruct dense 3D geometry for an observed scene. Based on the scene representations, traditional MVS methods can be mainly divided into three categories: voxel, point cloud, and depth map [9]. Voxel-based methods [10, 11, 12, 13] discretize the 3D space into voxels and label each as inside or outside of the surface. They are limited to small-scale scenes due to large memory consumption. Point cloud-based methods [14, 15] operate directly on 3D points and often employ propagation to gradually densify the reconstruction. As point cloud propagation occurs sequentially, these methods are difficult to parallelize, leading to longer processing times [9]. In addition, the irregularity and large size of point cloud are also not very suitable for deep learning, especially in large-scale scenes. In contrast, methods that are based on depth maps [16, 17, 18, 19, 20] use patch matching with photometric consistency to estimate the depth maps for individual images. Subsequently, these 2D depth maps are fused into a dense 3D representation, e.g., point cloud or mesh. This design decouples the reconstruction task into per-view depth estimation and depth fusion, which explicitly improves flexibility and scalability. In addition, storing a set of 2D depth maps to represent 3D geometry is more convenient [21, 17]. Although MVS has been studied extensively for several decades, traditional MVS methods rely on hand-crafted matching metrics and thus encounter challenges in handling various conditions, e.g., illumination changes, low-textured areas, and non-Lambertian surfaces [22, 23, 24, 25].
@@
In this survey, motivated by previous works [21, 15, 9], we categorize learning-based MVS methods based on their underlying representations as follows: depth map-based, voxel-based, NeRF-based, 3D Gaussian Splatting-based, and large feed-forward methods. Voxel-based methods estimate geometry with volumetric representation and Signed Distance Functions (SDF). They are limited to small-scale scenes because of the high memory storage of voxel grids. NeRF and 3D Gaussian Splatting-based methods adapt NeRF [28] and 3D Gaussian Splatting [29], which are used for novel view synthesis, and extract the surface from the volumetric implicit field and explicit 3D Gaussians points, respectively. They typically need to optimize the geometry for each scene. Large feed-forward methods typically use large models to directly learn the 3D representation from given images. They require massive computation because of the large network. Depth map-based methods introduce deep learning in depth estimation and then fuse depth maps with the traditional fusion algorithm. Similar to traditional MVS, depth map-based methods [9, 26] are the main family of learning-based MVS since they inherit the advantages of those traditional methods [30] by decoupling 3D reconstruction into depth estimation and depth fusion. Therefore, we focus more on depth map-based MVS in this survey.
For clarity, we further categorize depth map-based MVS into online and offline methods, shown in Fig. 1. Online methods [26, 31] reconstruct scenes in an online manner with local temporal information, while offline methods [9, 32] reconstruct scenes in an offline manner with global information. To estimate depth for a given reference view, online methods select its best views in a local buffer that stores previous views. In contrast, offline methods select best views among all possible views in the database.
In summary, our survey covers the most recent literature on learning-based MVS methods, including five main families: depth map-based, voxel-based, NeRF-based, 3D Gaussian Splatting-based, and large feed-forward methods. We provide a comprehensive review and insights on different aspects, including the pipelines and algorithmic intricacies. Moreover, we evaluate the performance and efficiency of the reviewed methods, and discuss the potential future directions for deep learning-based MVS.
@@ 2 Preliminaries
Depth map-based MVS, including most traditional and learning-based methods, typically consists of several components: camera calibration, view selection, multi-view depth estimation and depth fusion. In this section, we introduce these components to provide readers a clear picture of the MVS. Note that camera calibration and view selection are generally required by other methods as well. Sec. 2.1 introduces camera calibration with Structure from Motion (SfM) or SLAM. Sec. 2.2 discusses how to select neighboring views to reconstruct the geometry. Sec. 2.3 explains how to build cost volumes in learning-based MVS methods with plane sweep [33]. Sec. 2.4 introduces the typical depth fusion strategies after depth estimation. Sec. 2.5 lists common datasets and benchmarks for MVS and Sec. 2.6 summarizes common evaluation metrics.
[…]
MVS
Traditional Methods COLMAP [30], P-MVS [15], Gipuma [18] Datasets & Benchmarks ScanNet [27], DTU [22], Tanks and Temples [23] Learning-based MVS with Depth Representation Pipeline Camera Calibration, View Selection, Multi-View Depth Estimation, Depth Fusion Supervised
Online Methods
MVDepthNet [26], DeepvideoMVS [31], SimpleRecon [34]
Offline Methods
Direct 3D CNN MVSNet [9], CIDER [35] RNN R-MVSNet [36], $D^{2}$HC-RMVSNet [37] Coarse-to-fine CasMVSNet [32], UCSNet [38] Iterative Update PatchmatchNet [39], IterMVS [40] Unsupervised End-to-end JDACS [41], RC-MVSNet [42] Multi-stage U-MVSNet [43], KD-MVS [44] Semi-supervised SGT-MVSNet [45] Learning-based MVS with Other Representations Voxel Direct 3D CNN Atlas [46] Coarse-to-fine NeuralRecon [47], TransformerFusion [48] NeRF Optimization NeuS [49], Neuralangelo [50], UniSDF [51] Generalizable SparseNeuS [52], VolRecon [53], LRM [54] 3D Gaussian Splatting Optimization SuGaR [55], 2DGS [56], PGSR [57] Generalizable MVSplat [58], DepthSplat [59], LGM [60] Large Feed-forward Point-based DUSt3R [61], VGGT [62], MapAnything [63] Future Directions Dataset & Benchmarks, View Selection, Feature Extraction, Generative Reconstruction, Efficiency, Prior Assistance
@@ 2.1 Camera Calibration
Camera calibration is a process of determining the intrinsic and extrinsic parameters of a camera to understand its geometry and characteristics accurately [64]. It serves as the foundational step in MVS, ensuring that the subsequent reconstruction process is built on accurate and consistent geometric information, ultimately leading to a more reliable and precise 3D representation of the scene. Typically, obtaining calibrated camera parameters is usually achieved by running off-the-shelf SfM algorithms [17, 65] or SLAM [66], which jointly optimize sparse triangulated 3D points and camera parameters. The camera parameters include the extrinsic matrix $\mathbf{T}=[\mathbf{R}|\mathbf{t}]$ and intrinsic matrix $\mathbf{K}$. Depth map-based MVS methods [9, 32, 39] require a bounded depth range $d_{\textrm{min}},d_{\textrm{max}}]$ to improve the estimation accuracy. For offline methods [9, 32], the depth range can be estimated by projecting the sparse point cloud from SfM to each viewpoint and compute the minimum and maximum $z$ values [9]. In contrast, online methods [26, 31] usually set constant depth ranges, e.g., [0.25mm, 20.00mm], since the scene scale is usually fixed as room.
@@ 2.2 View Selection
First, for most online MVS methods [26, 67, 31, 34], a frame is selected as a keyframe when its pose has sufficient difference compared with the previous keyframe. Within a keyframe buffer, each new keyframe selects several old keyframes with smallest pose difference to estimate depth [31]. To measure this pose difference, GP-MVS [67] proposes a heuristic measure as:
[…]
where $\mathbf{T}_{ij}=[\mathbf{R}_{ij}|\mathbf{t}_{ij}]$ is the relative transformation between view $i$ and $j$. A new keyframe is added if the pose-distance $\text{disc}(\mathbf{T}_{ij})$ from the most recent keyframe is above a certain threshold [31, 34].
Second, for most offline MVS methods [9, 36, 32], view selection is done with the sparse point cloud obtained by SfM [30, 65]. For a reference view $i$, MVSNet [9] computes a score $s(i,j)=\sum_{\mathbf{P}}\eta(\mathbf{\Theta}_{ij}(\mathbf{P}))$ for the neighboring view $j$, where $\mathbf{P}$ is a 3D point observed by both view $i$ and $j$.
$\mathbf{\Theta}_{ij}(\mathbf{P})$ is the baseline angle for $\mathbf{P}$, which is computed as:
$\mathbf{\Theta}_{ij}(\mathbf{P})=(180/\pi)\arccos{((\mathbf{c}_{i})-\mathbf{P})\cdot((\mathbf{c}_{j})-\mathbf{P}))},$ (2)
where $\mathbf{c}_{i},\mathbf{c}_{j}$ are the camera centers. $\eta(\cdot)$ is a piece-wise Gaussian function [68] to favor a certain baseline angle $\theta_{0}$:
$\eta(\theta)=\left\{\begin{aligned} \exp{(-\frac{(\theta-\theta_{0})^{2}}{2\sigma_{1}^{2}})},\theta\leq\theta_{0}\\ \exp{(-\frac{(\theta-\theta_{0})^{2}}{2\sigma_{2}^{2}})},\theta>\theta_{0}\end{aligned},\right$. (3)
where $\theta_{0}$ is the favored baseline angle, and $\sigma_{0}$ and $\sigma_{1}$ are weights. Then the view selection is done by choosing neighboring views with highest scores. Almost all the following offline MVS methods [36, 32, 39] use the same strategy.
@@ 2.3 Multi-view Depth Estimation with Plane Sweep
To form a structured data format that is more suitable for convolution operations, most depth map-based MVS methods rely on the plane sweep algorithm [33] to calculate matching costs. The plane sweep algorithm discretizes the depth space with a set of fronto-parallel planes along the depth direction. This practice is deeply inspired by learning-based binocular stereo methods [69, 70, 71], which assess matching costs for a set of disparity hypotheses and subsequently estimate the disparity.
@@
In practice, we divide the depth range, which is manually set [26] or estimated by SfM [9], into discrete samples and assign hypotheses at these values. To map coordinates with depth hypothesis $d$, homography transformation [9] $\mathbf{p}_{i}(d)\sim\mathbf{H}_{i}(d)\cdot\mathbf{p}$, is applied, where $\mathbf{p}_{i}(d)$ is the corresponding pixel in the $i$-th source view for the pixel $\mathbf{p}$ of the reference view. The homography $\mathbf{H}_{i}(d)$ between the $i$-th source view and the reference view 0 can be computed as:
@@
where $\mathbf{K}_{0}$, $\mathbf{K}_{i}$ denote camera intrinsics, $\mathbf{R}_{0}|\mathbf{t}_{0}]$, $\mathbf{R}_{i}|\mathbf{t}_{i}]$ denote camera extrinsics, $\mathbf{n}$ denotes the principal axis of the reference view. Equivalently, we can also project a reference pixel in the source views with depth hypothesis [35, 39, 40]. We compute $\mathbf{p}_{i}(d)$ in the $i$-th source view for pixel $\mathbf{p}$ in the reference and depth $d$ as follows:
@@
The warped source feature is then obtained via differentiable bilinear interpolation of the source feature map. To better elaborate the process of plane sweep, we include a toy example in the supplement. The photometric similarity (or matching cost) is measured in a one-to-many manner between reference and warped source features for the following depth estimation, which will be discussed in Sec. 3.2.
@@ 2.4 Depth Fusion
For depth map-based MVS, after estimating all the depth maps, we need to fuse them into a dense 3D representation, e.g., point cloud or mesh. Online MVS methods [31, 34] usually adopt TSDF (Truncated Signed Distance Function) fusion [72, 6] to sequentially fuse depth maps into a TSDF volume and then use Marching Cube [73] to extract the mesh. However, there usually exist outliers in the depth maps, which may reduce the accuracy of reconstruction. To adress this problem, offline MVS methods [9, 32, 39] filter out outliers before converting the depth maps to point clouds. Depth filtering involves photometric and geometric filtering [9] (for more details, please refer to the supplement). Note that online MVS methods do not perform depth filtering before TSDF fusion. For depth fusion, offline MVS methods use the visibility-based fusion algorithm [74] to minimize depth occlusions and violations across different viewpoints. The fused depth maps are then directly reprojected to space to generate the 3D point cloud [9].
@@ 2.5 Datasets and Benchmarks
High-quality and large-scale 3D datasets and benchmarks are important for both training and evaluation. We briefly summarize the commonly used MVS datasets and benchmarks in  Tab. I.
SUN3D [75] is a RGB-D video dataset for large indoor scenes, which provides RGB-D images, camera poses, segmentations, and point clouds.
ScanNet [27] is a large RGB-D dataset that contains 1613 indoor scenes with ground-truth camera poses, depth maps, surface reconstruction, and semantic segmentation labels. Online MVS methods [26, 31, 34] mainly use ScanNet for training and testing.
7-Scenes [76] is a RGB-D dataset captured in indoor scenes with a handheld Kinect RGB-D camera. Since it is relatively small, 7-Scenes is usually used to test the generalization performance of the models trained on ScanNet [27] without finetuning [31].
DTU [22] is an object-centric MVS dataset collected under well-controlled laboratory conditions with known camera trajectory. It contains 128 scans with 49 or 64 views under 7 different lighting conditions. Since DTU dataset provides scanned ground truth point clouds instead of depth maps, it is required to generate mesh models with surface reconstruction, e.g., screened Poisson surface reconstruction algorithm [77], and then render depth maps [9] for training.
Tanks and Temples [23] is a large-scale benchmark captured in more complex real indoor and outdoor scenarios. It is divided into intermediate and advanced sets. Different scenes have different scales, surface reflection and exposure conditions. Note that Tanks and Temples does not provide ground truth camera parameters, which are usually estimated with Structure-from-Motion [17, 65].
ETH3D [24] contains 25 large-scale indoor and outdoor scenes with high-resolution RGB images. The scenes typically contain many low-textured regions and non-Lambertian surfaces, e.g., white walls and reflective floor. In addition, the images are usually sparse and have strong viewpoint variations and occlusions. Therefore, ETH3D is considered as a very challenging benchmark.
BlendedMVS [25] is a large-scale synthetic dataset for MVS training. It contains a variety of scenes, such as cities, sculptures, and shoes. The dataset consists of more than 17kk high-resolution images rendered with reconstructed models and is split into 106 training scenes and 7 validation scenes. Since images are rendered through virtual cameras, the camera parameters are accurate enough for training.
ScanNet++ [78] is a large-scale and high-resolution indoor dataset with 3D reconstructions, high-quality RGB images, commodity RGB-D video. Compared to ScanNet [27], the quality of color and geometry is higher.
TABLE I: An overview of commonly used datasets and benchmarks for learning-based Multi-View Stereo.
Dataset Year Num. Scene Complexity Provided Ground Truth Evaluation
Scenes Indoor Outdoor Camera Pose Depth Point Cloud Mesh Target
SUN3D [75] 2013 415 Depth
7-Scenes [76] 2013 7 Depth
DTU [22] 2016 124 Point Cloud
ScanNet [27] 2017 1503 Depth / Mesh
Tanks and Temples [23] 2017 21 Point Cloud1
ETH3D [24] 2017 25 Point Cloud1
BlendedMVS [25] 2020 113 Depth
ScanNet++ [78] 2023 460 Depth / Mesh
  • 1
    For datasets with online benchmark, the point cloud ground truth of test set is not released.